OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (mrnood):

first rewrite the equation so that it is in the form y= mx + b Post your rewritten equation here...

OpenStudy (anonymous):

tbh i got no clue. i already took 6 tests today and now i don't understand anything new anymore. but i have to finish 10 today

OpenStudy (mrnood):

3x - y = -3 If you cannot rewrite the equation then you need to study until you can - it is the only way to answer these questions. Follow these 2 steps in order then post your result here If you cannot follow these steps then you cannot answer the question... FIRST add 3 to BOTH sides of the equation SECOND ad y to BOTH sides of the equation That will give the equation in the form I wrote above. However, IF you are just going to wait for someone to give you the answer - it won't be me.....

OpenStudy (anonymous):

tbr i never asked for the answer. i just wanted someone to explain it to me in an easy way and not over 9 pages.

OpenStudy (mrnood):

OK - I have given you 2 simple steps above. That is not 9 pages. Do those steps and post your answer here

OpenStudy (anonymous):

the negative 3 or 3x??

OpenStudy (mrnood):

3x - y = -3 FIRST add 3 to BOTH sides of the equation SECOND ad y to BOTH sides of the equation

OpenStudy (anonymous):

3x-4y????

OpenStudy (mrnood):

that is not an equation - so you haven't done what I suggested. (and where did oyu get 4y from) Please just do the 2 simple steps exactly as suggested

OpenStudy (mrnood):

soz gtg

OpenStudy (anonymous):

x-y=-6?????? sorry but i am totally confused

OpenStudy (anonymous):

@SolomonZelman can you explain it to me a little more thoroughly so i know what and why im doin it?????

OpenStudy (anonymous):

@wolf1728 @firstyoufall @KlOwNlOvE @paki @funkypanda @Rogue